### v1

# Adaptive Text Watermark for Large Language Models

Abstract
The advancement of Large Language Models (LLMs) has led to increasing concerns about the misuse of AI-generated text, and watermarking for LLM-generated text has emerged as a potential solution. However, it is challenging to generate high-quality watermarked text while maintaining strong security, robustness, and the ability to detect watermarks without prior knowledge of the prompt or model. This paper proposes an adaptive watermarking strategy to address this problem. To improve the text quality and maintain robustness, we adaptively add watermarking to token distributions with high entropy measured using an auxiliary model and keep the low entropy token distributions untouched. For the sake of security and to further minimize the watermark’s impact on text quality, instead of using a fixed green/red list generated from a random secret key, which can be vulnerable to decryption and forgery, we adaptively scale up the output logits in proportion based on the semantic embedding of previously generated text using a well designed semantic mapping model. Our experiments involving various LLMs demonstrate that our approach achieves comparable robustness performance to existing watermark methods. Additionally, the text generated by our method has perplexity comparable to that of un-watermarked LLMs while maintaining security even under various attacks. Our code will be available at github.com/yepengliu/adaptive-watermark.
Machine Learning, ICML
1 Introduction
The rapid development of Large Language Models (LLMs) (Zhang et al., 2022; Touvron et al., 2023) has ushered in transformative possibilities but also brings forth potential challenges. One primary challenge lies in the misuse of AI-generated text for malicious purposes, such as spreading disinformation, generating fake news, or engaging in plagiarism. The sheer volume and quality of AI-generated text make it increasingly difficult to discern between human and AI-generated content. As a solution, implementing a watermark on LLMs proves to be an effective method for detecting AI-generated text.
Specifically, the watermarking technique encodes certain patterns into the text generated by LLMs, which are imperceptible to humans but can be detected by the algorithm. This is usually achieved by introducing well-designed perturbations to the original logits or using a special sampling method. Kirchenbauer et al. (2023a) design a watermark for LLM by randomly separating the vocabulary into ‘green’ and ‘red’ lists and increasing the frequency of tokens from the ‘green’ list. The previously generated token is used as a random key to determine the ‘green/red’ list for the next token. This method achieves good detection performance when the watermarked text remains unmodified. However, after modifications such as word replacement and paraphrasing, this watermarking method will fail as the ‘green/red’ list cannot be reliably recovered. As it is common for individuals to alter AI-generated text rather than use it directly, improving the robustness of the watermark becomes an essential requirement.
To improve the robustness of LLM watermarking, Zhao et al. (2023) propose to use a fixed ‘green/red’ list at each time step, which is generated from a random secret key that is shared with the detector. Both empirical and theoretical evidence demonstrate that this method substantially enhances robustness. More recently, Sadasivan et al. (2023) introduces a spoofing attack specifically designed to deduce hidden watermark signatures by analyzing the token frequency, thereby enabling the forgery of watermarked text. Such an attack offers a method for adversaries to circumvent the watermark, potentially damaging the reputation of the system. Regrettably, it has been shown that the watermarking algorithm proposed in Zhao et al. (2023) is vulnerable to spoofing attack, and one can infer the fixed ‘green’ token list and forge the watermarked text.
The studies above enhance the frequency of tokens in the ‘green’ list during generation by uniformly adding a fixed value to the logits of those ‘green’ tokens. This way of perturbing the distribution may change the distribution significantly, deteriorating the quality of the watermarked text. Therefore, Kuditipudi et al. (2023) presents a distortion-free watermarking method that maintains the original probability distribution of LLMs and embeds the watermark during the token sampling process. This method provides a watermarked text whose text perplexity is close to that of un-watermarked text. However, such a distortion-free watermarking lacks sufficient robustness against modifications, like paraphrasing.
Inspired by the aforementioned research, designing an effective watermark algorithm requires us to comprehensively consider at least the following three factors. (1) Robustness: The watermark should remain detectable after modifications without significantly changing the semantics of the watermarked text. (2) Security: The watermark should be resilient against efforts to reverse-engineer or decryption without authorization. (3) Text quality: LLMs-generated watermarked text should preserve a level of perplexity comparable to the un-watermarked text. Among existing works, achieving high-quality text often comes with the trade-off of robustness, and prioritizing strong robustness can potentially compromise security. Therefore, designing a watermark that simultaneously guarantees robustness, security, and text quality poses a significant challenge.
Figure 1: Workflow of the proposed Adaptive Watermark. Our Adaptive Watermark method will assess the entropy of the previously generated text at every time step and add watermarks only to high-entropy tokens. This figure illustrates two cases in the text generation process. For the first example, the next token distribution is high entropy as measured by AWTI, indicating high uncertainty. Then, the SLSVE module will extract the logits scaling vector based on the semantics of this text. Subsequently, we apply AWTS to perturb the sampling distribution. When it comes to the second text, it results in low entropy as measured by AWTI, suggesting a low uncertainty for the next token. In this situation, the next token will be directly sampled from the original probability distribution.
In this work, our goal is to provide a holistic approach that integrates robustness and security with the production of high-quality text. Additionally, we also ensure the watermark can be detected in an agnostic manner, meaning that the watermark detection process is independent of the original LLMs and prompts. Our watermark strategy consists of three steps as shown in Figure 1.
  • [leftmargin=*,topsep=0.1em,itemsep=-0.2em]
  • To minimize the impact of the watermark on the text quality while maintaining the robustness, we propose the Adaptive Watermark Token Identification (AWTI) so that the watermarking is added adaptively to high-entropy token distributions, and leaves the low-entropy token distributions intact. This method employs an auxiliary Measurement model to assess the entropy of each token, ensuring that the entropy can be reliably estimated during detection.
  • To further enhance the security of our method, we introduce the Semantic-based Logits Scaling Vector Extraction (SLSVE) to replace the widely used ‘green/red’ list. Our logits scaling vector is obtained by passing the semantics of the generated text (using a pre-trained Sentence embedding model) to a Semantic Mapping Model, which is hard to reverse-engineer.
  • We propose the Adaptive Watermark Temperature Scaling (AWTS), which proportionally scales up the original logits using the extracted logits scaling vector. Through this approach, the perturbation is applied by scaling the temperature of the distribution, which further reduces the influence on the text quality due to watermarking.
  • In the detection process, the AWTI is first applied to the text under analysis to identify potential watermarked tokens. After this, the SLSVE is employed to extract the logits scaling vectors and calculate the detection score.
We conduct extensive experiments over various language models (OPT-2.7B, OPT-6.7B (Zhang et al., 2022), GPT-J-6B (Wang & Komatsuzaki, 2021)) to show the effectiveness of our method. We employ perplexity as a metric to evaluate text quality, and our findings indicate that the quality of the watermarked text remains comparable to that of the un-watermarked text. To show the robustness of our method, we paraphrase the watermarked text using GPT3.5 and DIPPER (Krishna et al., 2023). The outcomes are on par with those achieved by Zhao et al. (2023) which is the most robust method in the literature. The spoofing attack is used to illustrate that our watermark is hard to decrypt. Lastly, our detection process is agnostic and does not rely on any original prompt and LLMs.
2 Related Work
As LLMs evolve, the need to distinguish AI-generated text from human-written text is crucial for authenticity, accountability, and intellectual property protection. Text watermarking, embedding hidden markers in text, emerges as a promising solution for this challenge. There are currently two main types of watermark methods in use: generated text watermarking and LLMs watermarking.
Generated text watermarking. Generated text watermarking primarily operates on existing text, altering aspects like format (Brassil et al., 1995; Por et al., 2012; Sato et al., 2023; Rizzo et al., 2016), lexical choices (Yoo et al., 2023; Yang et al., 2023; Munyer & Zhong, 2023; Yang et al., 2021), or syntax (Atallah et al., 2001; Topkara et al., 2006; Meral et al., 2009), while preserving the underlying semantics. EASYMARK (Sato et al., 2023), a format-based watermarking, utilizes subtle changes in Unicode characters and spacing to embed watermarks in text. For word-level modifications, Yang et al. (2021), based on context-aware lexical substitution, utilizes BERT to identify semantically suitable synonym candidates for watermarking. Meral et al. (2009) introduces a morphosyntax-based natural language watermarking scheme that uses syntactic tree diagrams to analyze the structure and functional dependencies of text. In addition to the aforementioned methods that primarily rely on specific rules and might lead to unnatural modifications, some neural-based approaches (Zhang et al., 2023; Abdelnabi & Fritz, 2021) involve encoding text with specific message signatures using a pre-trained language model. Abdelnabi & Fritz (2021) propose a method where both sentence embedding and message signatures embedding, extracted by an encoder, are fed into a decoder to generate the watermarked text
LLMs watermarking. Rather than adding a watermark to pre-existing text, an increasing number of studies are focusing on embedding watermarks directly into the text during the LLM generation process (Wang et al., 2023; Fairoze et al., 2023). Some research efforts are dedicated to enhancing the robustness of watermarks (Kirchenbauer et al., 2023a, b; Liu et al., 2023b; Zhao et al., 2023; Ren et al., 2023), aiming to ensure the persistence of watermarking even after exposure to various attacks. Zhao et al. (2023) introduce a provable robust watermark using a fixed vocabulary ‘green/red’ list generated using a shared random key. Ren et al. (2023) utilize the semantics of the generated text as a hash key to separate the vocabulary, which is achieved by discretizing the continuous embedding space. Some works focus on enhancing the security of watermarks (Hu et al., 2023; Wu et al., 2023; Christ et al., 2023; Liu et al., 2023a), aiming to make them difficult to forge. Hu et al. (2023) propose an unbiased watermark using two re-weighting methods. Liu et al. (2023b) proposed a semantic invariant watermark that balances robustness and security. However, it relies on original prompts for detection, which is often impractical to access. The watermarks mentioned above are inserted into the text by manipulating the logits produced by LLMs. Furthermore, some studies Kuditipudi et al. (2023); Christ et al. (2023) introduce watermarks using specially crafted sampling methods. These methods are unbiased and designed to ensure that the watermark has a minimal impact on the original LLM distributions to improve text quality.
Our work focuses on the LLMs watermarking. Previous works in this field have made significant strides in improving the robustness, text quality, and security aspects of watermarking techniques. However, it remains challenging to consider all these aspects simultaneously. Therefore, we intend to propose a holistic watermark approach that achieves strong robustness and high-quality text generation, while also being difficult to forge.
Notations. Given a Language Model (LM), denote $\mathcal{V}$ as the vocabulary of LM, and denote $\mathcal{|\mathcal{V}|}$ as the size of vocabulary. Given a input prompt $R=\{R^{(1)},R^{(2)},...,R^{(N)}\}$, the LM will generate a sequence of tokens $S=\{S^{(1)},S^{(2)},...,S^{(T)}\}$ based on the prompt. We denote $S_{0:t}=\{S^{(1)},...,S^{(t)}\}$ and $S_{0}=\emptyset$, and $r^{(n)}$ and $s^{(t)}$ are concrete tokens in the vocabulary, i.e., $r^{(n)}$, $s^{(t)}\in\mathcal{V}$. As an auto-regressive model, LM will predict the probability distribution $P_{LM}(\cdot|R,S_{0:t})$ over each token in $\mathcal{V}$ at time step $t+1$ based on the prompt and the preceding sequence, and sample the next token based on this probability distribution until the token of termination is sampled.
3 Adaptive Watermark Generation
In this section, we present our adaptive watermarking technique designed to minimize the impact on text quality, enhance watermark security, and maintain robustness simultaneously. Section 3.1 introduces Adaptive Watermark Token Identification to ensure watermark robustness while minimizing the impact of text quality in generated content. The Semantic-based Logits Scaling Vector Extraction, as discussed in Section 3.2, replaces the widely used ‘green/red’ list to separate the vocabulary, making it more difficult to forge. In Section 3.3, we propose a new distribution perturbation method to inject watermarking, further improving watermarked text quality.
3.1 Adaptive Watermark Token Identification
The Adaptive Watermark Token Identification (see, Figure 1) is designed to minimize the impact of the watermark on the text quality while maintaining its robustness. The watermarking is added in an adaptive manner, i.e., only the distribution of those tokens with high entropy will be perturbed by our watermark method.
Shannon Entropy  (Shannon, 1948) measures the uncertainty of a distribution, which is defined as $H(P_{X})=-\sum_{x\in\mathcal{X}}P_{X}(x){\log}P_{X}(x)$, where $P_{X}$ is a discrete probability distribution defined over the space $\mathcal{X}$, and $0\leq H(P_{X})\leq\log|\mathcal{X}|$. When the entropy is low, the distribution becomes more concentrated on specific tokens. Conversely, when the entropy is high, the distribution becomes more spread out with larger uncertainty. A high entropy distribution in LLMs means that the original distribution offers multiple reasonable choices during sampling. Thus, perturbing the high entropy distribution to insert certain patterns may only subtly impact the sampling process, ensuring that the sampled tokens are still reasonable in the context.
More specifically, given a prompt $R$, the LLM will generate a logits vector $l^{(t+1)}$ over all the tokens in $\mathcal{V}$ at time step $t+1$. The softmax function is then applied to convert the logits into a probability distribution $P_{LM}(\cdot|R,S_{0:t})$, where the probability of the $k$-th token in the $\mathcal{V}$ is $p^{(t+1)}_{k}={\rm exp}(l^{(t+1)}_{k})/\sum_{i=1}^{|\mathcal{V}|}{\rm exp}(l^{(t+1)}_{i}).$ The LLM will sample a token from this probability distribution as the next token using a sampling method, such as Top-$K$ sampling or Top-$p$ sampling. At different time steps, the randomness and uncertainty of the sampling distribution may vary based on the context of the generated text.
To ensure reliable entropy estimation during text generation and detection process without relying on the watermarked LLM, we use an auxiliary Measurement Model (MM) to compute the entropy of the generated text at each time step. The MM can be an auto-regressive model that is significantly smaller in scale compared to the watermarked LLMs, such as GPT-2 (Radford et al., 2019), OPT-350M (Zhang et al., 2022). The entropy measured with MM is expressed as
$H\left(P_{MM}(\cdot|S_{0:t})\right)=-\sum_{i=1}^{|\mathcal{V}|}p^{(t+1)}_{i}\log p^{(t+1)}_{i}.$ (1)
We set a predefined entropy threshold, denoted as $\alpha$, and avoid disturbing distributions whose entropy is below this threshold. Therefore, our watermark only perturbs those distributions with $H\left(P_{MM}(\cdot|S_{0:t})\right)\geq\alpha$. We refer to the token sampled from the intact distribution as un-watermarked token and the token sampled from the perturbed distribution as watermarked token. As the threshold $\alpha$ increases, two effects become evident. Firstly, there will be more un-watermarked tokens during the generation process. Secondly, the impact of watermarked tokens diminishes, mainly due to the higher inherent uncertainty in their high-entropy distributions.
In addition, due to the fact that only watermarked tokens are involved in the detection process as described in Section 4, adaptive watermarking also ensures robustness by filtering out un-watermarked tokens.
3.2 Semantic-based Logits Scaling Vector Extraction
The Semantic-based Logits Scaling Vector Extraction (see, Figure 1) is primarily developed to ensure the security of the watermark. The intuition of this approach is to make sure that the perturbation in the distribution is determined by the semantics of the generated text using neural networks so that the inserted watermarking cannot be easily decrypted by attackers.
Previous work uses the hash of the preceding token or a fixed random key to generate the ‘green’ and ‘red’ lists. However, such lists can be recovered by analyzing the token frequency in watermarked text using statistical approaches, e.g., spoofing attacks, which enable adversaries to forge watermarks. In contrast, the proposed SLSVE will adaptively generate the logits scaling vector to perturb the distribution based on the semantics of the generated text using a trainable neural network. The complex high-dimensional semantic space and neural networks further increase the complexity of the watermark, making it more difficult for adversaries to decrypt.
Algorithm 1 Semantic-based Logits Scaling Vector Extraction (SLSVE)
0:  Sentence Embedding Model (SE), Semantic Mapping Model (SM), Input sentence $S$.
1:  $u^{(t)}(S)={\rm SE}(S_{0:t-1})$
2:  $v^{(t)}(S)={\rm SM}(u^{(t)}(S);\theta)$
3:  $\hat{v}^{(t)}(S)=\big{[}\mathbbm{1}\{v_{1}^{(t)}(S)>0\},...,\mathbbm{1}\{v_{|\mathcal{V}|}^{(t)}(S)>0\}\big{]}$
3:  logits scaling vector $\hat{v}^{(t)}$.
In particular, for tokens with distributions satisfying $H(P_{MM}(\cdot|S_{0:t-1}))\geq\alpha$ at time step $t$, we proceed to extract logits scaling vector based on the semantics of generated text. First, we obtain the sentence embedding using a Sentence Embedding Model (SE), i.e., $u^{(t)}(S)={\rm SE}({S}_{0:t-1})$, where $u^{(t)}\in\mathbbm{R}^{L}$. We employ a pre-trained LM as the Sentence Embedding Model, e.g., Sentence-BERT (Reimers & Gurevych, 2019), to extract the sentence embedding. Then, we train a Semantic Mapping Model (SM) to convert the sentence embedding into a logits scaling vector, denoted as $v^{(t)}(S)={\rm SM}(u^{(t)}(S);\theta)$, where $v^{(t)}\in\mathbbm{R}^{|\mathcal{V}|}$ and $\theta$ denotes the parameters of SM.
Specifically, we train a feedforward neural network with multiple residual connections so that the logits scaling vector should satisfy the following properties.
(1) Smoothness with respect to semantics: As we need to recover the logits scaling vector to perform detection even with modifications in the watermarked text, SM needs to be insensitive to small changes in semantics to improve the robustness of the watermark. In order to make the logits scaling vector aligns with the change in semantic embedding, we want to make sure that if two different sentences $S$ and $S^{\prime}$ are close in embedding space, they will have similar logits scaling vector $v$ and $v^{\prime}$. We will minimize the loss $\big{|}D(u(S),u(S^{\prime}))-D(v(S),v(S^{\prime}))\big{|}$, where we use the Euclidean distance $D(a,b)=\sqrt{\sum_{i=1}(a_{i}-b_{i})^{2}}$ to measure the distance between two sentence embeddings.
(2) Uniform Perturbation: To make the watermark easy to detect, we want to ensure the logits scaling vector increases the probability of roughly half of the tokens in the vocabulary. For any sentence $S$ within the training dataset $\mathcal{T}$, we will ensure that $\sum_{i\in\mathcal{V}}\mathrm{sign}\big{(}v_{i}(S)\big{)}=0$. Therefore, an equal rate of positive and negative entries.
(3) Unbiased token preference: The logits scaling vector should be designed to prevent consistent bias towards specific tokens in the vocabulary. This ensures that specific tokens do not always exhibit a high probability in distributions across various semantic contexts, preventing their frequent recurrence in the watermarked text and improve security. Therefore, for any token $i$ in the vocabulary $\mathcal{V}$, we will ensure $\sum_{S\in\mathcal{T}}\mathrm{sign}\big{(}v_{i}(S)\big{)}=0$.
Therefore, considering all the above properties, the loss function of the Semantic Mapping Model is:
$\mathcal{L}(\mathcal{T},\theta)=$ $\sum_{S,S^{\prime}\in\mathcal{T}}\big{|}D(u(S),u(S^{\prime}))-D(v(S),v(S^{\prime}))\big{|}$ (2)
$+$ $\sum_{S\in\mathcal{T}}\big{|}\sum_{i\in|\mathcal{V}|}\mathrm{sign}(v_{i}(S))\big{|}+\sum_{i\in|\mathcal{V}|}\big{|}\sum_{S\in\mathcal{T}}\mathrm{sign}(v_{i}(S))\big{|}.$
After getting the logits scaling vector $(v^{(t)})$ using the trained model, we apply the indicator function to transfer every entry in $v^{(t)}$ to either 00 or $1$, i.e., $\hat{v}_{i}^{(t)}(S)=\mathbbm{1}\{v_{i}^{(t)}(S)>0\}$.
3.3 Adaptive Watermark Temperature Scaling
To further minimize the impact of perturbation on the original probability distribution, we introduce the Adaptive Watermark Temperature Scaling method (see, Figure 1). This technique enhances the consistency of the perturbation with the original distribution by adaptively adjusting the temperature of the distribution.
Specifically, we add watermark to the original LLM logits $l^{(t)}$ using temperature scaling:
$\hat{l}^{(t)}=l^{(t)}\cdot(1+\delta\cdot\hat{v}^{(t)}),$ (3)
where $\delta>0$ controls the strength of the watermark. In this way, we adaptively perturb the logit of each token by proportionally scaling up a factor of $(1+\delta\cdot\hat{v}^{(t)})$ using the obtained SLSVE. If the original logit $l_{i}^{(t)}$ is small, the $\hat{l}_{i}^{(t)}$ after the manipulation remains small, indicating that our watermark has a minimal effect on these logits. Our temperature scaling mainly affects tokens with large original logits $l_{i}^{(t)}$, causing the corresponding $\hat{l}_{i}^{(t)}$ to increase more significantly. In contrast to uniformly increasing the logits of tokens in the ‘green list’ as in previous work (Kirchenbauer et al., 2023a), our approach is to embed a watermark by modulating the temperature of the logits. Therefore, our watermark strength relies on the original distribution, which will make the generated watermarked text more coherent. As we will show in Figure 4, AWTS also contributes to the improvement of text quality.
Algorithm 2 Adaptive Watermark Generation
0:  Language model $P_{LM}$, measurement model $P_{MM}$, prompt $R$, watermark strength $\delta$, opening sentence $\hat{S}$, entropy threshold $\alpha$, measure threshold $M$.
1:  for $t=1,2,...,T$ do
2:     if $t\leq M$ then
3:        $\hat{v}^{(t)}=\mathrm{SLSVE}(\hat{S})$
4:        $q_{k}^{(t)}=\frac{\exp(l_{k}^{(t)}\cdot(1+\delta\cdot\hat{v}_{k}^{(t)}))}{\sum_{i\in\mathcal{V}}\exp(l_{i}^{(t)}\cdot(1+\delta\cdot\hat{v}_{i}^{(t)}))}$
5:        $s^{(t)}\sim q^{(t)}$.
6:     else
7:        if $H(P_{MM}(\cdot|S_{0:t-1})\geq\alpha$ then
8:           $\hat{v}^{(t)}=\mathrm{SLSVE}(S_{0:t-1})$
9:           $q_{k}^{(t)}=\frac{\exp(l_{k}^{(t)}\cdot(1+\delta\cdot\hat{v}_{k}^{(t)}))}{\sum_{i\in\mathcal{V}}\exp(l_{i}^{(t)}\cdot(1+\delta\cdot\hat{v}_{i}^{(t)}))}$
10:           $s^{(t)}\sim q^{(t)}$.
11:        else
12:           $p_{k}^{(t)}=\frac{\exp(l_{k}^{(t)})}{\sum_{i\in\mathcal{V}}\exp(l_{i}^{(t)})}$,
13:           $s^{(t)}\sim p^{(t)}$.
14:        end if
15:     end if
16:  end for
16:  watermarked sequence $s^{(1)},s^{(2)},...,s^{(T)}$.
Algorithm 2 describes the whole process of our Adaptive Watermark algorithm. At the beginning of the text generation process, only a limited number of tokens can be used for extracting semantic embeddings, which makes the detection of the watermark unstable under attack. To maintain the stability and robustness of the watermark, we set a secret opening sentence $\hat{S}$. When the number of generated tokens is smaller than a pre-determined measure threshold $M$, we extract the semantic embedding from the opening sentence and apply SLSVE.
4 Adaptive Watermark Detection
In this paper, we propose to approximate the Likelihood Ratio Test (LRT) to detect watermark. Consider the following null hypothesis $H_{0}$ and alternative hypothesis $H_{1}$,
$H_{0}:{\rm The\ candidate\ text\ is\ not\ watermarked.}$
$H_{1}:{\rm The\ candidate\ text\ is\ watermarked.}$
Given a sequence of tokens $S=\{S^{(1)},S^{(2)},...,S^{(T)}\}$, we input the sequence to MM to calculate the entropy and select the potential watermarked token at each time step, following the same process as the AWTI described in Section 3.1. Denote the set of potential watermarked token as $\mathcal{W}$, i.e., for all $t\in\mathcal{W}$, $H(P_{MM}(\cdot|S_{0:t-1}))\geq\alpha$ holds. For any $t\in\mathcal{W}$, if it is sampled from the perturbed distribution, the distribution would be $P_{H_{1}}(\cdot|R,S_{0:t})$. Conversely, if it is sampled from the original distribution, the distribution would be $P_{H_{0}}(\cdot|R,S_{0:t})$. Subsequently, the logits scaling vector is extracted utilizing both the Semantic Embedding Model and the Semantic Mapping Model, as detailed in Section 3.2. As we perturb the distribution with temperature scaling, the likelihood ratio can be written as
${\log}~{}\frac{P_{H_{1}}(S^{(t)}=k|R,S_{0:t-1})}{P_{H_{0}}(S^{(t)}=k|R,S_{0:t-1})}$
$=~{}{\log}~{}\frac{e^{l_{k}\cdot(1+\delta\cdot\hat{v}_{k})}}{\sum_{i}e^{l_{i}\cdot(1+\delta\cdot\hat{v}_{i})}}+{\log}~{}\frac{e^{l_{k}}}{\sum_{i}e^{l_{i}}}$
$=~{}l_{k}\cdot(\delta\cdot\hat{v}_{k})+{\log}~{}\frac{\sum_{i}e^{l_{i}}}{\sum_{i}e^{l_{i}\cdot(1+\delta\cdot\hat{v}_{i})}}$
$\approx~{}l_{k}\cdot(\delta\cdot\hat{v}_{k}),$ (4)
where we omitted the superscript of time $t$ and the normalization factor in the softmax function.
Figure 2: Workflow of Adaptive Detection. This figure illustrates a single time step during the detection process. At each time step, we will first check if the current token is the potential watermarked token by applying AWTI to assess the entropy of the current distribution. For potential watermarked token, the preceding text will be used to extract the logits scaling vector through SLSVE. If the corresponding value of the potential watermarked token in the logits scaling vector is positive, it will be added to the total score.
In the detection process, the original logits $l_{k}$ are unknown, and our focus is only on identifying which tokens in the logits have been scaled up. For this purpose, we simply use $\delta\cdot\hat{v}_{k}$ as the test score to facilitate detection. Therefore, the test score is given by
Score(${\rm Score}(S=\{s^{(1)},...,s^{(T)}\})=\frac{\sum_{t\in W}\delta\cdot\hat{v}^{(t)}_{s^{(t)}}}{|W|},$ (5)
where $|W|$ represents the number of all potential watermarked tokens.
5 Experiment
5.1 Experiment Setting
Dataset and Prompt. The realnewslike subset in C4 dataset (Raffel et al., 2020) is used to test the performance of our methods. We use the first two sentences in each text as a prompt for LLMs, and the subsequent $200$ tokens as the human-generated text. We also generate $200\pm 30$ tokens using the LLMs and prompts. The Semantic Mapping Model is trained using the Sentence-Compression dataset  [^1] https://huggingface.co/datasets/embedding-data/sentence-compression.
Baseline and Language Model. We apply our method to three different models, which are OPT-2.7B, OPT-6.7B (Zhang et al., 2022) and GPT-J-6B (Wang & Komatsuzaki, 2021). At each time step, the next token is sampled from the distribution using a combination of Top-$K$ and Top-$p$ sampling methods, with $K=50$ and $p=0.9$. In our watermarking algorithm, the default hyperparameters are set as follows: $\alpha=2$, $\delta=1.5$, $M=50$. We use Sentence-Transformers (Reimers & Gurevych, 2019) to extract the sentence embedding. The specific model we used is all-mpnet-base-v2  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Our method is compared with existing methods, including KGW-00 (Zhao et al., 2023), KGW-$1$/KGW-$2$ (Kirchenbauer et al., 2023a), EXP-edit (Kuditipudi et al., 2023). In KWG-$1$/KWG-$2$, the numeral denotes the hashing of the previous 1 or 2 tokens, respectively. For KGW-00, KGW-$1$, and KGW-$2$, the ‘green list’ token percentage is set to be $0.5$.
Evaluation Methods. Watermark detection performance is evaluated through ROC-AUC value and the Best F1 Score. In addition, We present the ROC curve, which illustrates the true positive rate (TPR) and false positive rate (FPR) at various thresholds. The TPR and FPR indicate the accurate identification of watermarked text and the incorrect recognition of human-written text as being watermarked. The perplexity of the text is calculated using LLama-13B (Touvron et al., 2023), measuring the quality of the text.
5.2 Robustness
To demonstrate the robustness of our watermark strategy, we execute a paraphrase attack on the watermarked text using two different paraphrasing methods: GPT-$3.5$ and DIPPER (Krishna et al., 2023). For GPT-$3.5$, we use the gpt-3.5-turbo-instruct version, and we employ the same prompt as used in Kirchenbauer et al. (2023b) to paraphrase the watermarked text. DIPPER is a paraphrase generation model to rewrite the text and evade the watermark detection, and the lexical diversity is set to be $60$.
The results are summarized in Table 1, which shows that all methods demonstrate effective performance when the watermarked texts remain unaltered. However, when the watermarked text is paraphrased using GPT-$3.5$ and DIPPER, for OPT-$6.7$, the ROC-AUCs of our method are $0.981$ and $0.964$, respectively. Additionally, the best F1 score achieves $0.932$ and $0.903$. This performance significantly exceeds all the compared baselines. The experimental results on OPT-2.7 and GPT-J-6B also show that our watermark method is more robust than other methods under paraphrase attack.
Table 1: Performance evaluation of watermarked text under different conditions, including cases with no attack and paraphrase attacks.
Language Model Setting ROC-AUC Best F1Score
KGW-0 KGW-1 KGW-2 EXP-edit Ours KGW-0 KGW-1 KGW-2 EXP-edit Ours
OPT-$2.7$ No Attack $0.996$ $0.998$ $0.998$ $0.997$ 1.000 $0.989$ $0.997$ $0.999$ $0.994$ 1.000
GPT-3.5 $0.971$ $0.881$ $0.768$ $0.919$ 0.983 $0.916$ $0.818$ $0.719$ $0.871$ 0.952
DIPPER $0.939$ $0.846$ $0.767$ $0.899$ 0.960 $0.879$ $0.783$ $0.726$ $0.841$ 0.902
OPT-$6.7$ No Attack $0.995$ $0.995$ 0.999 $0.993$ $0.996$ $0.997$ $0.997$ $0.998$ $0.995$ 0.998
GPT-3.5 $0.964$ $0.874$ $0.769$ $0.881$ 0.981 $0.904$ $0.815$ $0.728$ $0.801$ 0.932
DIPPER $0.936$ $0.842$ $0.795$ $0.883$ 0.964 $0.867$ $0.782$ $0.743$ $0.807$ 0.903
GPT-J-6B No Attack 1.000 $0.997$ $1.000$ $0.997$ $0.998$ $0.995$ $0.993$ 1.000 $0.995$ $0.996$
GPT-3.5 $0.946$ $0.874$ $0.764$ $0.931$ 0.966 $0.887$ $0.804$ $0.721$ $0.868$ 0.913
DIPPER $0.923$ $0.850$ $0.803$ $0.960$ 0.963 $0.855$ $0.783$ $0.753$ $0.898$ 0.901
5.3 Text Quality
The quality of the generated text is evaluated by perplexity. A lower perplexity value suggests better text quality. As shown in Figure 3, we compare the perplexity of text generated by our method with human-written text, un-watermarked text (texts generated by the model without any watermarking), and other baseline methods. Our method has a negligible impact on the perplexity compared to the un-watermarked text. It also shows comparable results with EXP-edit, which is a distortion-free watermark method. The quality of text watermarked using our approach is clearly superior to that of KGW-00 and KGW-$1$.
Figure 3: Comparison of text perplexity among human-written text, un-watermarked text, and texts using various watermark methods across different language models. For KGW-0 and KGW-1, the watermark strength and green list size are set as $2.0$ and $0.5$, respectively.
5.4 Security Results
The spoofing attack (Sadasivan et al., 2023) is used to evaluate the security of our method, which is an adversarial attack designed to produce non-AI text that can be mistakenly identified as AI-generated by the watermark detector. This is accomplished by frequently querying the watermarked language model and examining the frequency of the 181 most commonly used words within the vocabulary. For our method, we use the sentences in the DBPedia Classes Dataset  [^3] https://www.kaggle.com/datasets/danofer/dbpedia-classes/data as prompt to generate watermarked text and analyze the word frequency. DBPedia Classes Dataset has three levels of hierarchy (L1, L2, L3). From L1 to L3, the granularity of classification becomes increasingly detailed, and the range of contexts within a specific class narrows down for different sentences. We compare the decryption rate among different watermark methods, which is the percentage of ‘green’ list tokens that the attacker can decrypt. From Table 2, we can observe that KGW-0 exhibits the highest decryption rate because of its fixed ’green/red’ list. For KGW-1 and KGW-2, the implementation of the previous token as a hash key adds a certain level of difficulty to the decryption process compared to the KGW-0. Our method demonstrates superior resistance to attack compared to those works, with a decryption rate of only $0.571$.
Table 2: Security performance for different watermark methods.
Language Model Evaluation KGW-0 KGW-1 KGW-2 Ours (L3)
OPT-6.7 Decryption Rate $0.963$ $0.912$ $0.723$ $0.571$
5.5 Ablation Study
In this section, our investigation focuses on assessing the robustness and text quality of watermarked text using our watermark under varying entropy thresholds and watermark strength. Specifically, for this analysis, we adjust the threshold value $\alpha$ to be 0, 1, 2, 3 and set the watermark strength $\delta$ at levels of $0.5$, $1$, $1.5$, and $2$.
Entropy Thresholds. In Table 3, we provide the ROC-AUC, Best F1 Score under different entropy thresholds when paraphrasing is conducted using GPT-3.5. In addition, we provide the Average Watermark Rate (AWR), which is the fraction of the watermarked tokens. Figure 4 shows the text perplexity comparison under different entropy thresholds. When $\alpha=0$, AWTI is skipped in watermark generation. The ROC-AUC and Best F1 Score are high, but the watermarked text exhibits high perplexity, with a median value exceeding $10$. When setting $\alpha$ to 1, 2, or 3, there is a notable decrease in the perplexity of the watermarked text, as fewer tokens are perturbed reflected by the decreasing AWR. In addition, the robustness of the watermark is improved when AWTI is applied ($\alpha=1,2$), potentially because some modified tokens are filtered out by AWTI in detection. For entropy threshold $\alpha=3$, our method can achieve results comparable to other methods, with only about half of the watermarked tokens.
Table 3: Robustness performance for different entropy thresholds.
Language Model Evaluation $\alpha=0$ $\alpha=1$ $\alpha=2$ $\alpha=3$
OPT-6.7 ROC-AUC $0.978$ $0.988$ $0.981$ $0.972$
Best F1 Score $0.935$ $0.957$ $0.932$ $0.917$
AWR $1.000$ $0.835$ $0.707$ $0.553$
Figure 4: Comparison of text perplexity at varying entropy thresholds and across different watermark strengths. $\alpha$ represents the entropy threshold. $\delta$ represents the watermark strength.
Watermark Strength. In Table 4 and Figure 4, we increase the watermark strength $\delta$ from $0.5$ to $2$ to examine its impact on robustness and perplexity. As shown in Table 4, there is little impact on the ROC-AUC values in terms of robustness. Interestingly, in contrast to previous works (Kirchenbauer et al., 2023a; Zhao et al., 2023) where perplexity tends to rise with larger watermark strength, the perplexity of the text watermarked with our method decreases as the $\delta$ increases in a reasonable range. This benefits from the proposed AWTS, which embeds the watermark by adjusting temperature as explained in Section 3.3. However, this does not imply that a larger $\delta$ consistently results in better text quality. A higher watermark strength can cause the token selection process to resemble greedy sampling more closely, potentially leading to the generation of repetitive content.
Table 4: Robustness performance for different watermark strength.
Language Model Evaluation $\delta=0.5$ $\delta=1$ $\delta=1.5$ $\delta=2$
OPT-6.7 ROC-AUC $0.982$ $0.984$ $0.981$ $0.978$
Best F1 Score $0.938$ $0.945$ $0.932$ $0.938$
6 Conclusion
In this paper, we propose a holistic LLM watermark method to achieve the following properties: strong robustness, high security, high text quality, and accurate detection without the need to access the original language model and prompt. Our experiments show that, compared to baseline methods, our approach demonstrates better performance under different paraphrase attacks. Notably, our method has a negligible impact on text perplexity when compared to un-watermarked text. Additionally, we highlight the security of our approach by employing the spoofing attack. Specifically, the capacity of language models differs among various models. Therefore, the model owner may need to tune the entropy threshold and watermark strength carefully based on the performance of the language model to get better watermarking performance. Moreover, the watermarked model owners are suggested to train their own small-sized Measurement Model based on the watermarked model for performance and security considerations. In future research, we aim to conduct a theoretical analysis to deepen the understanding of the performance characteristics of the watermarking approach and offer an optimal solution.
Appendix A Additional Experiment Results
A.1 Semantic Mapping Model Training Details
The Semantic Mapping Model is a neural network that starts and ends with linear layers and contains two residual blocks in the middle. For the loss function provided in Section 3.2, the weight of each term is set as 1. For the training process, the optimization algorithm we utilize is stochastic gradient descent.
When we train the Semantic Mapping Model (detailed in Section 3.2), in order to ensure the logits scaling vector is sufficiently resilient to variations of semantics, we rescale the Euclidean distance between two sentence embeddings $d=D(u(S),u(S^{\prime}))$ to a wider range. The original range $d\in[L,U]$, where $L=0$ and $U$ is the maximum distance. To enlarge this range, we apply a linear transformation $(T:\mathbb{R}\rightarrow\mathbb{R})$ to the original distance $d$, resulting in a new distance $d^{\prime}=T(d)\in[L^{\prime},U^{\prime}]$ where $L^{\prime}<L$ and $U^{\prime}>U$. The linear transformation is expressed as
$T(d)=\left(\frac{d-L}{U-L}\right)\cdot(U^{\prime}-L^{\prime})+L^{\prime},$
where we set the $L^{\prime}=-2$ and $U^{\prime}=4$ in our experiment.
A.2 Additional Robustness Results
Figure 5 illustrates the ROC curves and the corresponding AUC values when text is watermarked by different watermarking methods under paraphrasing attack. It is evident that our watermark method exhibits better robustness against paraphrase attacks compared to the other watermarking methods.
(a) Paraphrasing attack using GPT-3.5.
(b) Paraphrasing attack using DIPPER.
Figure 5: Comparisons of ROC curves of different watermark methods applied to various language models, assessing their performance against two different paraphrasing attacks.
A.3 Additional Ablation Study
Measurement Models. We investigate how different Measurement Models (MMs) impact the robustness of our watermarks and the quality of the generated text. For this purpose, we employ MMs with different sizes and types, such as GPT2 (124M), GPT2-large (774M), OPT-125M, and OPT-350M, in our evaluation. For all models, we set the $\alpha$ and $\delta$ as $2$ and $3$, respectively. As shown in Table 5, there is only a minor influence on the robustness of the watermark when different MMs are adopted.
However, the effect on text quality is more nuanced. As indicated in Figure 6, models like GPT2-large and OPT-350M will result in lower perplexity compared to the GPT2 and OPT-125M. One primary reason is that different MMs produce different entropy estimates, which results in different AWRs even with the same entropy threshold. Generally, a higher AWR implies a larger number of watermarked tokens, which can marginally impact text quality.
Table 5: Robustness performance for different Measurement Models.
Language Model Evaluation GPT2 GPT2-large OPT-125M OPT-350M
OPT-6.7 ROC-AUC $0.985$ $0.981$ $0.987$ $0.982$
Best F1 Score $0.951$ $0.932$ $0.947$ $0.938$
AWR $0.778$ $0.707$ $0.755$ $0.718$
Figure 6: Text perplexity comparison between different Measurement Models.
Effectiveness of Each Component in Adaptive Watermark. Our proposed watermarking method mainly consists of three key components: Adaptive Watermark Token Identification (AWTI), Semantic-based Logits Scaling Vector Extraction (SLSVE), and Adaptive Watermark Temperature Scaling (AWTS). Here, we use KGW-0 as our baseline model and analyze the robustness and perplexity of watermarked text by sequentially adding our three components. Table 6 and Figure 7 exhibit the robustness and text perplexity performance under the following four different settings. KGW-0+AWTI represents applying AWTI to the KGW-0, aiming to adaptively perturb the distributions with low entropy. KGW-0+AWTI+SLSVE indicates that both AWTI and SLSVE are applied to KGW-0. KGW-0+AWTI+SLSVE+AWTS represents our complete Adaptive Watermark, where all our three components are applied.
Table 6 and Figure 7 demonstrate that the application of AWTI reduces the perplexity of the watermarked text without compromising robustness performance. The introduction of SLSVE enhances the security of the watermark but comes at the cost of increased perplexity in the watermarked text. Finally, when all three components are applied, our watermark method excels in both robustness and perplexity performance compared to existing baselines.
Table 6: Robustness performance comparison after the removal of different components.
Language Model Settings ROC-AUC Best F1 Score
OPT-6.7 KGW-0 $0.964$ $0.904$
KGW-0 + AWTI $0.962$ $0.904$
KGW-0 + AWTI + SLSVE $0.960$ $0.893$
KGW-0 + AWTI + SLSVE + AWTS $0.981$ $0.932$
Figure 7: Text perplexity comparison after the removal of different components.
A.4 Security Analysis Details
In this subsection, we provide a detailed introduction to the security analysis outlined in Section 5.4. The security of a watermark primarily relies on the difficulty of decryption by potential attackers. This paper uses the spoofing attack (Sadasivan et al., 2023) to decrypt the watermarking scheme. The spoofing attack aims to uncover the potential ‘green’ token of the watermarking scheme by analyzing the 181 common token frequencies in the watermarked text.
For KGW-0, which uses a fixed ‘green/red’ token list throughout the generation, we directly calculate the frequency of tokens in the watermarked text. In the case of KGW-1/KGW-2, the previous one/two token is used as a hash key to generate a ‘green’ list. Therefore, we examine the occurrences of token pairs in the watermarked text to ascertain the ‘green’ list corresponding to each specific token. In our method, the logits scaling vector varies according to the semantics. We use L3 level data (Musical class) in the DBPedia Classes Dataset as the prompt to generate the watermarked text. For all watermarking methods, we identify the top 50 highest frequency words from the 181 common tokens and calculate the percentage of these words that are actually ‘green’ tokens. For each watermarking method, we produce 5000 instances of watermarked texts, each consisting of 200 tokens.
Appendix B Watermarked Text Generation time
This section compares the watermarked text generation time and the watermark detection time for different approaches. For each method, we generate and detect 100 instances of watermarked text, each containing roughly 200 tokens. Subsequently, we compute the average time taken for both the generation and detection of each instance of watermarked text. For our method, we set the value of the entropy threshold $\alpha=2$ and the watermark strength $\delta=1.5$. We evaluate the generation and detection time of the proposed adaptive watermarking with two different Measurement models.
Table 7 presents all the results. For generation time, despite employing three additional models in our watermarked text generation process, our average generation time remains faster than that of KGW-0 and KGW-1. Our method is slightly slower than EXP-edit, primarily because EXP-edit uses a sampling-based watermarking approach that does not require logit computation. Specifically, using a Measurement model with smaller parameters can lead to faster generation times. Regarding the time required for detection, our method is slightly slower than KGW-0 and KGW-1. However, it is significantly faster than EXP-edit, which necessitates computing the alignment between the watermarked text and the watermark key sequence.
Table 7: Watermarked text generation and detection time across various watermarking methods.
Language Model Watermarking Method Settings Average Generation Time Average Detection Time
OPT-6.7 KGW-0 - $81.64$s $0.01$s
KGW-1 - $85.15$s $1.73$s
EXP-edit - $25.89$s $163.19$s
Ours GPT2-large $33.16$s $5.57$s
Ours OPT-125M $31.15$s $3.63$
Appendix C Watermarked Text Example
We present some examples of watermarked text using our proposed method. Table 8 shows the un-watermarked text and watermarked text under different entropy thresholds ($\alpha$). The watermarked texts are generated using the OPT-6.7B model with the watermark strength $\delta=1.5$. The watermarked tokens are highlighted in yellow. As the value of $\alpha$ increases, it is evident that the proportion of watermarked tokens noticeably decreases.
Table 8: Examples of watermarked text using our method with various entropy thresholds. In these examples, tokens highlighted in yellow represent watermarked tokens, indicating that these tokens were sampled from a perturbed distribution.
Prompt
It’s not always easy to keep mentally focused during running. Focus requires discipline and a certain level of mental toughness, meaning that you have developed the capacity to perform to the best of your ability despite external conditions or internal distractions.
Un-watermarked text
While it’s not impossible to train for a specific event and keep mentally focused at all times, it’s challenging. Here are three things you can do to help yourself.
Plan Ahead
The best time to think about race strategy is during warm-ups before any race. Once your legs get moving, they will tell you what they need—you just have to listen to them! It’s also a good idea to write down in detail the training plan for your event and make sure that you understand it completely. Plan your warm-up so that there is an emphasis on building speed without injury risk.
Get Out of Your Head
There are lots of other things to focus your attention on besides running. If you’re running around town, take a walk break every quarter mile or so. This allows you to stretch, cool off and rest, while keeping your muscles engaged.
Watermarked text ($\alpha=1$)
The ability and desire for discipline are developed through practice. The first step in developing discipline during running, especially if a runner wants a higher level performance, begins with setting goals and working towards meeting these goals. Goal setting allows runners an avenue to create the mental toughness needed to achieve their best results during training runs or races. Setting realistic expectations will also allow runners to work within their abilities instead of trying to perform beyond what they are currently capable of. Setting realistic expectations can also serve as motivation for runners to continue training despite the obstacles that life has to offer during the course of training or race day. It can also serve as motivation to continue training even when injuries or other problems make training challenging or downright unpleasant. For runners who want to set realistic expectations for themselves during training or race day, the below checklist can serve as guidelines:
Mentally prepared - Have enough sleep the night before the race or training runs
Stress free environment - Remove negative people from the environment
Eat right - Eat healthy meals the day before the race or training runs
Watermarked text ($\alpha=2$)
The best runners are often described in terms like, focused, determined, and driven. You may have read that somewhere on a running magazine cover, but did anyone actually write that down? I was recently asked if I knew any specific exercises to help people stay mentally focused during training runs. My response wasNo,” but the question got me thinking more about how to develop mental toughness for runners. Staying mentally prepared is an important part of the training process that can be just as important as physical preparation, but many runners don’t give it much thought. It’s easy to see how physical preparation is important for race day performance, but what exactly does it mean when you talk about mental preparation? Here are five things you need to do to stay mentally prepared while running:
1. Have realistic goals and expectations. Some people set unrealistic expectations for themselves or others regarding the number of miles they’ll run each week, their times for certain races, or even their overall goal for their training program.
Watermarked text ($\alpha=3$)
The best runners are focused on a specific task, which allows for a level playing field. The same applies in any aspect in which concentration and discipline are necessary. In order for a runner, cyclist, tennis star, golffer, etc. to perform their best, they must have the ability to stay focused at all times. To achieve this state of focus or mental toughness requires training (mental conditioning) and practice. You do not have to run every day for hours, but just like anything else it’s about putting in the time and practicing until you get better over time.
So how does this apply to your life? The best runners know that running is an important part of their lives, as well as their training routine and their overall health. The best runners also understand what is required to keep them mentally strong during the race. They know how to train themselves so that they will always be ready and prepared for the next race, regardless of external conditions or the obstacles within themselves.